The Magellanic Clouds Newsletter

An electronic exchange of Magellanic Clouds information

Edited by: You-Hua Chu, Dominik J. Bomans, and Eva K. Grebel

                                 mcnews@astro.uiuc.edu
                      http://www.astro.uiuc.edu/mcnews/MCNews.html
                 http://www.astro.uni-wuerzburg.de/~grebel/MCNews.html

No. 11 June 2, 1997

Table of Contents

News and Views  1
Abstracts of 14 refereed papers  2
Abstracts of 4 non-refereed papers 10
Thesis Abstract  12

News and Views

A Proposed Meeting on the Magellanic Clouds

A proposal has been submitted to the IAU to hold a symposium on “New Views of the Magellanic Clouds” in Victoria, Canada, on July 13-17, 1998. Further information about the proposal can be found at http://www.astro.uiuc.edu/mcnews/MC symp proposal.html and http://www.astro.uni-wuerzburg.de/˜ grebel/MC symp proposal.html. Comments and suggestions can be sent to mcnews@astro.uiuc.edu for the time being.

Four IAU Divisions (IV, V, VI, and VIII) and ten IAU Commissions (25, 27, 28, 29, 33, 34, 45, 47, 40, and 44) agree to support this proposed symposium. The proposal will be reviewed by the IAU Executive Committee; their decisions will be made in August 1997 at the Kyoto meeting. Please express your support for the proposed symposium to the IAU Executive Committee before then. Thank you.

Abstracts of Refereed Papers

H_{2} and its relation to CO in the LMC and other Magellanic irregular galaxies

F.P. Israel^{1}

1 Sterrewacht Leiden, Postbus 9513, 2300 RA Leiden, the Netherlands

H_{2} column densities towards CO clouds in the LMC and SMC are estimated from their far-infrared surface brightness and HI column density. The newly derived H_{2} column densities imply N(H_{2})/I(CO) conversion factors (in units of 10^{21} mol cm^{−2} ( K km s^{−1})^{−1}) X_{LMC} = 1.3±0.2 and X_{SMC} = 12±2. LMC and SMC contain total (warm) H_{2} masses of 1.0±0.3 × 10^{8} M_{⊙} and 0.75±0.25 × 10^{8} M_{⊙} respectively. Local H_{2}/HI mass ratios similar to those in LMC and SMC are found in the Magellanic irregulars NGC 55, 1569, 4214, 4449 and 6822 and in the extragalactic HII region complexes NGC 604, 595 and 5461 in M 33 and M 101 respectively. In these HII regions and in NGC 4449, we find X = 1–2; in NGC 55, 4214 and 6822 X = 3–6 again in units of 10^{21} mol cm^{−2} ( K km s^{−1})^{−1}. The post-starburst galaxy NGC 1569 has a very high value similar to that of the SMC.

The CO–H_{2} conversion factor X is found to depend on both the ambient radiation field intensity per nucleon σ_{FIR}/ N_{H} and metallicity [O]/[H]: log X ∝ 0.9±0.1 log \frac{σ_{FIR}}{N_{H}} - 3.5±0.2 log\frac{[O]}{[H]}

Neglecting the dependence on radiation field, our sample also yields a reasonable approximation: log X ∝ -2.7±0.3 log\frac{[O]}{[H]}

Milky Way values are consistent with these relations. This result is interpreted as the consequence of selective photodissociation of CO subjected to high radiation field energy densities and poor (self)shielding in low-metallicity environments, and especially the preferential destruction of diffuse CO in ‘interclump’ gas.

Although locally H_{2} may be the dominant ISM-component, the average global H_{2}/HI mass ratio is 0.2±0.04 and the average H_{2} gas mass fraction is 0.12±0.02. Magellanic irregulars have warm molecular gas fractions very similar to those of our Galaxy, whereas other global properties (mass, luminosity, metallicity, CO luminosity) are very different.

Accepted by: Astronomy & Astrophysics

For preprints, contact israel@strw.leidenuniv.nl

Dust in the Small Magellanic Cloud: Interstellar Polarization and Extinction

C. V. Rodrigues^{1,2}, A. M. Magalhães^{1}, G. V. Coyne, S. J.^{3} and V. Piirola^{4}

1 Instituto Astronômico e Geofísico - USP, Caixa Postal 9638, 01065-970 São Paulo - SP, Brazil 2 Inst. Nacional de Pesquisas Espaciais, Div. Astrofísica, Caixa Postal 515, 12201-970 São José dos Campos - SP, Brazil 3 Vatican Observatory, V-00120 Vatican City State, Rome, Italy 4 Tuorla Observatory - University of Turku, SF-21500 Piikkiö, Finland

The typical extinction curve for the Small Magellanic Cloud (SMC), in contrast to that for the Galaxy, has no bump at 2175 Å and has a steeper rise into the far ultraviolet. For the Galaxy the interpretation of the extinction and, therefore, the dust content of the interstellar medium has been greatly assisted by measurements of the wavelength dependence of the polarization. To the present no such measurements existed for the SMC.

Therefore, to further elucidate the dust properties in the SMC we have for the first time measured linear polarization in five colors in the optical region of the spectrum for a sample of reddened stars. For two of these stars, for which there were no existing UV spectrophotometric measurements, but for which we measured a relatively large polarization, we have also obtained data from the International Ultraviolet Explorer (IUE) in order to study the extinction. With the help of parameterization, we attempted to correlate the SMC extinction and polarization data. In addition, we performed dust model fits to both extinction and polarization using silicate and graphite or amorphous carbon spheres and silicate cylinders. The size distribution for the cylinders is taken from a fit to the polarization and we introduce the notion of volume continuity between this and the silicate sphere size distribution.

The main results are: (1) the wavelength of maximum polarization, λ_{max}, in the SMC is typically smaller than that in the Galaxy; (2) however, AZV 456, which shows the UV extinction bump, has a λ_{max} typical of that in the Galaxy, its polarization curve is narrower, its bump is shifted to shorter wavelengths as compared to the Galaxy and its UV extinction does not conform to the Galactic analytical interpolation curve based on the ratio of total to selective extinction; (3) the ’typical’, monotonic SMC extinction curve can be best fit with amorphous carbon and silicate grains; (4) the extinction towards AZV456 may only be explained by assuming a larger gas-to-dust ratio than the observed N(HI)/A(V) value, with a small amount of the available carbon in graphite form; (5) from an analysis of both the extinction and polarization data and our model fits it appears that the SMC has typically smaller grains than those in the Galaxy.

The absence of the extinction bump in the SMC has generally been thought to imply a lower amount or even an absence of carbon in solid form in the SMC compared to the Galaxy. Our results show that the size distribution, and not only the carbon abundance, is different in the SMC as compared to the Galaxy. In addition, and contrary to previous findings, another component besides silicates is indeed needed to provide a sizeable part of the observed extinction towards the SMC. Using the SMC as a laboratory for studying the solid component of the interstellar medium, we also discuss some of the implications of our results in view of proposed interstellar dust models.

Accepted by : The Astrophysical Journal For preprints, contact claudia@das.inpe.br

Also available by anonymous ftp at ftp://crux.das.inpe.br/pub/claudia/smc

Kinematics of the very young nebula N59 at the edge of the supershell LMC4

M. Rosado^{1}, A. Laval^{2}, E. Le Coarer^{3}, J. Boulesteix^{2}, P. Amram^{2}, A. Viale^{2}

1 Instituto de Astronomía, UNAM, Apartado Postal 70-264, CP04510, México, D.F. 2 Observatoire de Marseille, 2 Place Le Verrier, 13248 Marseille, Cedex 4, France 3 Observatoire de Grenoble, Domaine de St. Martin d’Heres, France

The dynamics of the nebula N59 , located at the boundary of the supershell LMC4, in the Large Magellanic Cloud , is studied using a scanning Fabry-Perot interferometer. It is shown that the nebulae NGC 2032 and 2035, which form the bright core of the HII region N59A, are a single HII region which looks separated due to the presence of a heavy dust lane. This bright core presents an expansion motion of 24 km s^{−1}. The kinetic energy involved in this motion is of about 1.5 x 10 ^{49} erg. This value is compatible both with a supernova explosion origin or with a formation by the winds of interior massive stars. Since no clear traces of a SN explosion have been found in this nebula and since the stellar content of N59A is rich in blue stars, we conclude that these stars, mainly the very massive star HDE 269810 (R122), and probably other stars hidden by the dust lane, are largely sufficient to provide the wind power driving the expansion motion. The dust lane seems to be mixed up with the nebular gas and the stars, suggesting a site where star formation may still take place.

An extended shell, probably ionized by the star R122, has been detected at the same velocities as the blueshifted slab which is seen in absorption at the foreground of LMC4. The star R122 contributes to the high excitation of the faint diffuse gas, and perhaps of some farther nebulae.

To the East, the SNR 0536-676 remains as a trace of the existence of another massive star that already exploded. The kinematics of N59B, which contains the SNR 0536-676, is also studied, corroborating the results of previous studies.

Submitted to : Astronomy & Astrophysics

For preprints, contact margarit@astroscu.unam.mx

Two Ring Nebulae around Blue Supergiants in the Large Magellanic Cloud

K. Weis^{1,2}, Y.-H. Chu^{2}, W.J. Duschl^{1,3}, and D.J. Bomans^{2}

1 Institut für Theoretische Astrophysik, Tiergartenstr. 15, D-69121 Heidelberg, Germany 2 University of Illinois, Department of Astronomy, 1002 W. Green Street, Urbana, IL 61801, USA 3 Max-Planck-Institut für Radioastronomie, Auf dem Hügel 69, D-53121 Bonn, Germany

Ring nebulae are often found around massive stars such as Wolf-Rayet stars, OB and Of stars and Luminous Blue Variables (LBVs). In this paper we report on two ring nebulae around blue supergiants in the Large Magellanic Cloud. The star Sk−69 279 is classified as O9f and is surrounded by a closed shell with a diameter of 4.5 pc. Our echelle observations show an expansion velocity of 14 km s^{−1} and a high [Nii]λ6583Å/Hα ratio. This line ratio suggests nitrogen abundance enhancement consistent with those seen in ejectas from LBVs. Thus the ring nebula around Sk−69 279 is a circumstellar bubble.

The star Sk−69 271, a B2 supergiant, is surrounded by an Hα arc resembling a half shell. Echelle observations show a large expanding shell with the arc being part of the approaching surface. The expansion velocity is ∼ 24 km s^{−1} and the [Nii]λ6583Å/Hα is not much higher than that of the background emission. The lack of nitrogen abundance anomaly suggests that the expanding shell is an interstellar bubble with a dynamic age of 2 × 10^{5} yr.

Accepted by : Astronomy & Astrophysics

For preprints, contact kweis@ita.uni-heidelberg.de

Also available from the URL

   http://www.ita.uni-heidelberg.de/publications/preprints/1997/index.html

Obscured Asymptotic Giant Branch stars in the Magellanic Clouds III. New IRAS counterparts

Jacco Th. van Loon^{1,2}, Albert A. Zijlstra^{1}, Patricia A. Whitelock^{3}, L.B.F.M. Waters^{2},

Cecile Loup^{4,5}, Norman R. Trams^{6}

1 ESO-Garching, Germany, ^{2} Amsterdam, The Netherlands, ^{3} SAAO, South Africa, ^{4} ESO-Santiago, Chile, ^{5} IAP, France, ^{6} ESA-Vilspa, Spain

We have searched for near-infrared stellar counterparts of IRAS point sources in the Large Magellanic Cloud (LMC), in J and K-bands. This resulted in the detection of 21 counterparts, of which 19 are new discoveries. Using colour-magnitude and colour-colour diagrams, we identify 13 Asymptotic Giant Branch (AGB) stars with thick circumstellar dust envelopes, 7 possible early post-AGB stars or stars recovering from a thermal pulse, and 1 red supergiant or foreground star. For 10 of the IRAS targets we do not succeed in detecting and/or identifying a near-infrared counterpart. We serendipitously detect 14 other red sources, of which 2 are known Long Period Variables, and a few galaxies. The near-infrared and optical colours of the galaxies may indicate considerable interstellar extinction through the LMC, as much as A_{V} about 2-4 mag. The relative number of AGB carbon stars over oxygen stars is shown to decrease as the luminosity increases. Yet amongst the faintest mass-losing AGB stars oxygen-rich stars still exist, which puts constraints on current convection theories that predict the occurrence of third dredge-up and Hot Bottom Burning. We investigate the nature of some LMC stars that have infrared properties very similar to suspected Galactic post-AGB stars.

Accepted by : Astronomy & Astrophysics

Preprints available from the URL http://xxx.lanl.gov/abs/astro-ph/9704201

Cepheids in NGC 1866 - a Test for Pulsational Models

G. Bono^{1} and M. Marconi^{2}

1 Trieste Astronomical Observatory, Via G. B. Tiepolo 11, 34131 Trieste, Italy 2 Dept. of Physics, University of Pisa, Piazza Torricelli 2, 56100 Pisa, Italy

We present nonlinear theoretical results concerning the predicted pulsation behaviour of stellar models suitable for the sample of Cepheids belonging to the Large Magellanic Cloud (LMC) star cluster NGC 1866. The blue and red edges of the instability strip, transformed into the observational plane (V, B-V) by adopting current values for the distance modulus (DM=18.57 mag., Welch et al. 1991) and for the reddening (E(B-V)=0.06 mag., Arp 1967; Walker 1974; Walker 1987), are in agreement with the observed distribution of Cepheids in the quoted cluster. Moreover, the distribution of observational data in the Bailey diagram (V amplitude versus period) appears in agreement with theoretical predictions derived adopting the luminosity level predicted by canonical evolutionary models. The role played by the direct integration of the coupling between pulsation and convection and by the input physics on the smoothness of light and velocity curves is briefly discussed.

Accepted by : Monthly Notices of the Royal Astronomical Society

For preprints, contact marcella@astr1pi.difi.unipi.it

Chemical composition of six K supergiants in the SMC

V. Hill^{1}

1Observatoire de Paris, Section de Meudon, DASGAL, URA 335 du CNRS, 92195 Meudon Cedex, France

A sample of six SMC cool stars (K supergiants) are analysed by high resolution spectroscopy in order to investigate their chemical content, which should reflect that of the current interstellar medium within the SMC. The abundance of Na, Al, α-elements, Fe peak and heavy elements are derived through an LTE analysis of the lines, compared to a Galactic “standard” (the well-studied K giant Arcturus) and discussed in terms of chemical evolution of the SMC. The following results were obtained: 1.) All the stars from this sample are metal deficient ([Fe/H] ranges from –0.59 to –0.89 dex). The abundance varies little from star to star (0.10 dex r.m.s., comparable with the expected error due to uncertainties on the effective parameters of the stars), in agreement with the very small chemical composition variations among the H ii regions within the SMC found by Pagel et al. 1978. 2.)

Sodium is not strongly enhanced in these stars, and the star-to-star scatter is larger than expected. 3.) In contrast to metal-poor stars in our Galaxy, the so-called α-elements do not seem to be enhanced. 4.) Nickel displays a noticeable depletion with respect to iron ([Ni/Fe]≈ −0.3 dex), at variance with what is found in the SMC F supergiants and in Arcturus. 5.) Except for PMMR 144, the s and r process elements heavier than Ba are enhanced in all of our stars by ≃ +0.4 dex.

Accepted by : Astronomy & Astrophysics

For preprints, contact Vanessa.Hill@obspm.fr

Carbon, nitrogen, oxygen, and lithium abundances of six cool supergiants in the SMC

V. Hill^{1}, B. Barbuy^{2}, M. Spite^{1}

1Observatoire de Paris, Section de Meudon, DASGAL, URA 335 du CNRS, 92195 Meudon Cedex, France 2Universidade de São Paulo, IAG, Dept. de Astronomia, CP 9638, São Paulo 01065-970, Brazil

Carbon, nitrogen and oxygen abundances were derived from high-resolution spectra of 6 cool supergiants of the Small Magellanic Cloud.

Oxygen-to-iron ratios (mean value \overline{[O/Fe]} ≈–0.18 dex) are found to be similar to those found in young objects in the LMC and the Galaxy. This result is discussed in terms of chemical evolution. A mean deficiency of the carbon-to-iron ratio of \overline{[C/Fe]} ≈–0.3 dex and a nitrogen-to-iron ratio \overline{[N/Fe]} ≈+0.22 dex might imply a mixing signature, which is confirmed by the low ^{12}C/^{13}C ratio measured in three stars. In terms of chemical evolution, carbon plus nitrogen abundances have to be considered: in our sample, a mean value of \overline{[(C + N)/Fe]} ≈ –0.15 dex indicates a slight deficiency in carbon plus nitrogen with respect to the Sun, similar to the deficiencies found in Galactic supergiants and Orion. Lithium has been detected in all the program stars, reaching ϵ(Li)=0.6 dex for two stars.

Accepted by : Astronomy & Astrophysics

For preprints, contact Vanessa.Hill@obspm.fr

HST Observations of the Field Star Population in the Large Magellanic Cloud

Rebecca A. W. Elson^{1}, Gerard F. Gilmore^{1} and Basilio X. Santiago^{2}

1 Institute of Astronomy, Madingley Rd.,Cambridge CB3 0HA, UK 2 Federal Rio Grande do Sul, Universidade, Instituto de Física, Brazil

We present V and I photometry, obtained with the Hubble Space Telescope, for ∼ 15, 800 stars in a field in the inner disk of the Large Magellanic Cloud. We confirm previous results indicating that an intense star formation event, probably corresponding to the formation of the LMC disk, occurred a few times 10^{9} years ago. We find a small but real difference between our field and one further out in the disk observed by Gallagher et al. (1996): either star formation in the inner disk commenced slightly earlier, or the stars are slightly more metal rich. We also find evidence for a later burst, around 1 Gyr ago, which may correspond to the formation of the LMC bar. About 5% of the stars in our field are substantially older than either burst, and are probably members of an old disk or halo population with age ∼ 12 Gyr.

Accepted by : Monthly Notices of the Royal Astronomical Society For preprints, contact elson@ast.cam.ac.uk

The LMC transition star R 84 and the core of the LH 39 OB association

M. Heydari-Malayeri^{1}, F. Courbin^{2}, G. Rauw^{2}, O. Esslinger^{3}, and P. Magain^{2}

1 demirm, Observatoire de Paris, 61 Avenue de l’Observatoire, F-75014 Paris, France 2 Institut d’Astrophysique, Université de Liège, 5, Avenue de Cointe, B-4000 Liège, Belgium 3 Department of Physics and Astronomy, University of Wales, College of Cardiff, CF2 3YB, Cardiff, UK

On the basis of sub-arcsecond imaging obtained at the ESO NTT with SUSI and the ESO ADONIS adaptive optics system at the 3.6 m telescope, we resolve and study the core components of the LMC OB association LH 39. The central star of the association, the rare transition object R 84, is also investigated using CASPEC echelle spectroscopy at the ESO 3.6 m telescope. A new, powerful image restoration code that conserves the fluxes allows us to obtain the magnitudes and colors of the components. We bring out some 30 stars in a ∼ 16′′×16′′ area centered on R 84. At a resolution of 0′′.19 (fwhm), the closest components to R 84 are shown to be stars #21 and #7 lying at 1′′.1 NW and 1′′.7 NW respectively of the transition star. The former is possibly a blue star of V = 16.7 mag and the latter with its V = 17.5 mag is the reddest star of the field, after R 84. Star #7 turns out to be too faint to correspond to the red M2 supergiant previously reported to contaminate the spectrum of R 84. If the late-type spectrum is due to a line-of-sight supergiant with a luminosity comparable to R 84, it should lie closer than 0′′.12 to R 84. The transition star shows spectral variability between 1982 and 1991. We also note some slight radial velocity variations of the Of emission lines over time scales of several years. Furthermore, we derive the spectral types of two of the brightest stars of the cluster, using long slit spectra obtained at the NTT telescope equipped with EMMI, and discuss the apparent absence of O type stars in this association.

Accepted by: Astronomy & Astrophysics

For preprints, contact heydari@mesioa.obspm.fr

The Discovery of 8.9 Second Pulsations from the Variable X–Ray Source 2E0050.1–7247 in the Small Magellanic Cloud

G.L. Israel^{1}, L. Stella^{1}, L. Angelini^{2}, N. E. White^{2}, P. Giommi^{3} and S. Covino^{4}

1 Osservatorio Astronomico di Roma, V. dell’Osservatorio 2, I–00040 Monteporzio Catone (Roma), Italy 2 Laboratory for High Energy Astrophysics, Code 662, NASA–Goddard Space Flight Center, Greenbelt, MD 20771, USA 3 SAX Science Data Center, ASI, Viale Regina Margherita 202, I–00198 Roma, Italy 4 Osservatorio Astronomico di Brera, Via E. Bianchi 46, I–22055 Merate (Lecco), Italy

During a systematic search for periodic signals in a sample of ROSAT PSPC light curves, we discovered ∼ 8.9 s X–ray pulsations in 2E0050.1–7247, a variable X–ray source in the Small Magellanic Cloud. The source was detected several times between 1979 and 1993 at luminosity levels ranging from ∼ 5 × 10^{34} erg s^{−1} up to 1.4 × 10^{36} erg s^{−1} with both the Einstein IPC and ROSAT PSPC. The X–ray energy spectrum is consistent with a power law spectrum which steepens as the source luminosity decreases. We revealed a pronounced Hα activity from at least two B stars in the X–ray error circles. These results strongly suggest that the X–ray pulsar 2E0050.1–7247 is in a Be–type massive binary.

Accepted by : Astrophysical Journal Letters

For preprints, contact gianluca@coma.mporzio.astro.it
Also available from the URL http://www.mporzio.astro.it/~gianluca/publi.html/

Modelling the temporary X-ray off-state of CAL 83

P. Kahabka

Astronomical Institute and Center for High Energy Astrophysics, University of Amsterdam, Kruislaan 403, 1098 SJ Amsterdam, The Netherlands

The X-ray light curve of CAL 83 as measured with the ROSAT HRI covering ∼ 1.5 years (from July 95 to Dec 96) is derived. These observations show the pre-off state (persistent state), the April 96 off-state and the post off-state which returns after some decline to its previous persistent level. The intensity variations are discussed in the model of variations of the temperature of the envelope of the nuclear burning white dwarf. In order to explain the observed X-ray off-state a drop in temperature from 4.5 × 10^{5} K to at least 3.2 × 10^{5} K (i.e. by 40%) is required. This might have been achieved - assuming an accretion rate close to Ṁ_{RG} of a ∼ 1M_{⊙} white dwarf - if the radius of the envelope of the white dwarf increased by a factor of ∼2.0 possibly due to an increased mass accretion episode onto the white dwarf. CAL 83 resembles then the supposed wide binary supersoft source (and symbiotic star) AG Dra where a radius expansion by a factor of ∼2 and a temperature drop by ∼35% has been required in order to model the observed X-ray fall in 1994. In both systems the accretion rate is supposed to be close to Ṁ_{RG} and to pass episodically above this value and to cause the X-ray fall. The main difference between the two systems are the nature of the donor star, the mass of the white dwarf, and the size of the binary orbit. It is not known whether CAL 83 is similar to the transient supersoft X-ray transient RX J0513.9-6951 as no recurrent X-ray off-states have been found for CAL 83. The extreme short duration of < 100 days of the X-ray off-state together with the fact that the white dwarf is massive (M_{W D} > 1.1 M_{⊙}) constrains an X-ray recurrence time to ∼1-5 years. But possibly CAL 83 is not consistent with a conventional recurrent supersoft source. The X-ray off-state in CAL 83 may have been triggered by an enhanced mass accretion event.

Submitted to : Astronomy & Astrophysics For preprints, contact ptk@astro.uva.nl

Upper Limit to the Distance to the Large Magellanic Cloud

A. Gould^{1} and O. Uza^{1}

1 Dept of Astronomy, Ohio State University, Columbus, OH 43210, USA

We remeasure the “light echo” times to the near and far side of the ring around SN 1987A using the method of Gould (1995) and the new reductions of the original UV spectra by Sonneborn et al. (1996). We obtain an upper limit to the distance to the Large Magellanic Cloud μ_{LMC} < 18.37 ± 0.04. This result is very close to Gould’s original value, and contradicts claims that the new spectra measurements lead to a significant increase in the distance estimate and a substantial increase in the statistical errors. We elaborate on Gould’s original argument that the upper limit is robust, i.e., that it is only weakly sensitive to unmodeled aspects of the ring such as its geometry, hydrodynamic properties, or ionization history.

Submitted to : The Astrophysical Journal

For preprints, contact gould@astronomy.ohio-state.edu

Microlensing of tidal debris on the Magellanic great circle

H. Zhao^{1,2}

1 Sterrewacht Leiden, Niels Bohrweg 2, 2333 CA, Leiden, The Netherlands (present address) 2 Max-Planck-Institut fr̈ Astrophysik, Karl-Schwarzschild-Straße 1, 85740 Garching, Germany

Increasing evidence suggests that the Galactic halo is lumpy on kpc scales due to the accretion of at least a dozen small galaxies (LMC/SMC, Sgr, Fornax etc.). Faint stars in such lumpy structures can significantly microlens a background star with an optical depth 10^{−7} − 10^{−6}, which is comparable to the observed value to the LMC (Alcock et al. 1996).

The observed several microlensing events towards the LMC can be explained by a tidal debris tail lying in the Magellanic plane (Lynden-Bell 1976) if the progenitor of the Magellanic Clouds and Stream and other satellite galaxies in the Magellanic plane has a mass about twice that of the disk of the LMC. The LMC stars can either lens stars in the debris tail behind the LMC, or be lensed by stars in the part of debris tail in front of the LMC.

The models are consistent with an elementary particle dominated Galactic halo without massive compact halo objects (MACHOs). They also differ from Sahu’s (1994) LMC-self-lensing model by predicting a higher optical depth and event rate and less concentration of events to the LMC center. The debris stellar populations are detectable with good color-magnitude diagrams and standard candles in directions towards the Magellanic Clouds.

Submitted to: Monthly Notices of the Royal Astronomical Society

For preprints, contact hsz@strw.LeidenUniv.nl
Also available from the URL http://babbage.sissa.it/abs/astro-ph/9703097

Abstracts of Non-Refereed Papers

SN 1987A - Presupernova Evolution and the Progenitor Star

S.E. Woosley^{1}, A. Heger^{2}, T.A. Weaver^{3}, and N. Langer^{4}

1 Board of Studies in Astronomy and Astrophysics, University of California, Santa Cruz, CA 95064, USA 2 Max Planck Institut für Astrophysik, D–85740 Garching, Germany 3 General Studies Group, Physics Department, Lawrence Livermore National Laboratory, Livermore, CA 94550, USA 4 Institut für Theoretische Physik und Astrophysik, Universität Potsdam, D–14415 Potsdam, Germany

Ten years later, astronomers are still puzzled by the stellar evolution that produced SN 1987A — a blue supergiant. In single star models, the new OPAL opacities make blue solutions more difficult to achieve, though still possible for certain choices of convection physics. We also consider rotation, which has the desirable effect of producing large surface enhancements of nitrogen and helium, but the undesirable effect of increasing the helium-core mass at the expense of the envelope. The latter makes blue solutions more difficult. Still, we seek a model that occurs with high probability in the LMC and for which the time-scale for making the last transition from red to blue, ∼20 000 years, has a physical interpretation — the Kelvin-Helmholtz time of the helium core. Single star models satisfy both criteria and might yet prove to be the correct explanation for Sk -69 202, provided new rotational or convection physics can simultaneously give a blue star and explain the ring structure. Some speculations on how this might be achieved are presented and some aspects of binary models briefly discussed.

To appear in: SN 1987A: Ten years after, eds. M.M. Phillips and N.B. Suntzeff, PASPC

For preprints, contact ahg@MPA-Garching.MPG.De

Also available from the URL http://www.mpa-garching.mpg.de/{ ahg/Publications/Preprints/1987A-10.ps.gz

Fossils in the LMC: Measuring the Star Formation History from Field Stars

K. Olsen^{1}, P. Hodge^{1}, A. Dolphin^{1}, N. Suntzeff^{2}, and A. Walker^{2}

1 Astronomy Dept., Box 351580, University of Washington, Seattle, WA 98195, USA 2 Cerro Tololo Inter-American Observatory, NOAO, Casilla 603, La Serena, Chile

The star formation history of the Large Magellanic Cloud is recorded in its numerous field stars. By examining this fossil record, one can study the different roles external influences and stochastic processes play in global star formation. Through analysis of the distribution of stars in color-magnitude diagrams (CMDs) of the field, it is possible to recover the parameters that determine the star formation history. We present V − I, V color-magnitude diagrams of fields near six old LMC clusters obtained with WFPC2. These CMDs extend in magnitude from the upper main sequence and red giant branch to the unevolved lower main sequence. Our strategy for deducing the star formation history from the CMDs will involve comparing the observed distribution of stars with predictions based on model isochrones and artificial star tests. By using an automated technique to compare the observations with the predictions, we intend to arrive at best-fit solutions to the star formation rate and metallicity as functions of time. Previous studies of LMC field stars have revealed that the galaxy appears to have undergone a burst of star formation ∼ 2 − 4 billion years ago. We aim to build on the previous studies by improving the age resolution with which the star formation history is measured and exploring whether there is evidence for a spatial variation in the star formation history, as might be expected if external interactions trigger bursts of star formation.

Presented at: Northwest Astronomers Meeting, Yakima, WA, May 16-17 1997

The HD 5980 binary system: Components and Spectral Types

Virpi S. Niemela^{1}, Rodolfo H. Barbá^{1}, Nidia I. Morrell^{1}, and Mariela Corti^{1}

1 Facultad de Ciencias Astronómicas y Geofísicas, Universidad Nacional de La Plata, 1900 La Plata, Argentina

We present a radial velocity study of the optical spectral lines of the eclipsing binary HD 5980 in the Small Magellanic Cloud, which recently underwent a major luminous blue variable (LBV)-type outburst. From our data we confirm that both binary components appear as stars with emission line spectra of Wolf-Rayet type. The binary component which is in front of the system during the primary eclipse, is the more massive one, and the one which produced the outburst. Absorption lines present in the pre-outburst spectrum of HD 5980 seem not to follow the orbital motion of the binary, but show variations from epoch to epoch.

To appear in : Luminous Blue Variables: Massive Stars in Transition (eds. A. Nota & H. Lamers), ASP Conf. Ser.

For preprint, contact virpi@fcaglp.fcaglp.unlp.edu.ar
Also available by anonymous ftp at ftp://163.10.4.47/pub/virpi/HD5980a.ps
or from the URL http://sol.stsci.edu/~rbarba/full papers/kona hd5980b.ps.gz

Evolution of the Optical Spectrum and Hydrogen Abundances in HD 5980

Rodolfo H. Barbá^{1}, Virpi S. Niemela^{1}, and Nidia I. Morrell^{1}

1 Facultad de Ciencias Astronómicas y Geofísicas, Universidad Nacional de La Plata, 1900 La Plata, Argentina

We review the several types of spectral variations observed in the optical spectrum of HD 5980, the eclipsing binary in the Small Magellanic Cloud that unexpectedly underwent a luminous blue variable (LBV)-type outburst in 1994. We have also determined the Hydrogen presence in the spectrum, and find a correlation between the spectral type exhibited by the star along the outburst cycle, and the deduced amount of Hydrogen.

To appear in : Luminous Blue Variables: Massive Stars in Transition (eds. A. Nota & H. Lamers), ASP Conf. Ser.

For preprints, contact virpi@fcaglp.fcaglp.unlp.edu.ar
Also available by anonymous ftp at ftp://163.10.4.47/pub/virpi/HD5980b.ps
or from the URL http://sol.stsci.edu/~rbarba/full papers/kona hd5980a.ps.gz

Thesis Abstracts

Chemical Evolution of the Magellanic Clouds: – a study of the young stellar population in the Small and the Large Cloud – an approach to the old population of the Large Cloud

V. Hill^{1}

1 Observatoire de Paris, Section de Meudon, DASGAL, URA 335 du CNRS, 92195 Meudon Cedex, France

The Magellanic Clouds (MCs) are the two closest galaxies: these irregular dwarf satellites of our Galaxy, in addition to their outstanding role in calibrating various distance indicators, offer moreover a unique opportunity to test stellar and galactic theories. This work is dedicated to the study of the chemical evolution of the Magellanic Clouds, through the analysis of the detailled abundances of young stars: a sample of nine Large Magellanic Clouds (LMC) F supergiant stars, and of six Small Magellanic Cloud (SMC) K supergiants. Among the results discussed in this work, the following may be emphasised: 1- No abundance gradients are observed in either the Large or the Small Cloud. 2- the analysis of C and N shows that products of the CNO cycle have been dredged up towards the surface of the star, which is further confirmed by the carbon isotopic ratio ^{12}C/^{13}C (first measurement in the SMC). 3- The lithium abundance in the SMC is found to vary from one star to the other, suggesting different efficiencies for the mixing. 4- The O/Fe and a to iron ratios are comparable in the supergiants of the SMC, the LMC and our Galaxy, the simplest explanation being a parallel evolution of the three galaxies, and a milder star formation in the MCs. 5- The (C+N) abundance and the C/O ratio seem to be different from those found in the H ii regions. 6- The ratio of very heavy elements (produced through the s- and r- processes: La, Eu) to the heavy elements (Y, Sr), is higher than observed in our Galaxy for similar objects.

On the other hand, we show how the knowledge of the abundances of older populations in the MCs would help discriminating between various star formation processes (occurring continuously or in short bursts). In this respect, we present a new approach to the old population (Pop II) of the LMC, through the analysis of several thousand RR Lyrae type variable stars, recently detected by EROS (an experiment searching for dark matter through gravitational microlensing). We showed in particular, how the mean metal deficiency of this population challenges our current knowledge of the star formation history in the LMC.

Ph.D. Thesis

For offprints, contact Vanessa.Hill@obspm.fr